OpenStudy (anonymous):

find the greatest common factor of following monomials pq5 pq2

OpenStudy (cggurumanjunath):

PQ^5 AND PQ^2

OpenStudy (cggurumanjunath):

PQ^5 = P * Q * Q * Q * Q * Q PQ^2 = P * Q * Q GCD = P * Q * Q = P Q^2

OpenStudy (cggurumanjunath):

@waymansims1

OpenStudy (cggurumanjunath):

is that clear ?

OpenStudy (cggurumanjunath):